OpenStudy (anonymous):

rename 1/7 and 3/5 using the least common denominator

OpenStudy (callmekiki):

You're adding them?

OpenStudy (anonymous):

Least common denominator. What does 5 and 7 go into equally?

OpenStudy (anonymous):

no i just need to rename them

OpenStudy (callmekiki):

LCD: 7 times 5

OpenStudy (callmekiki):

What is that?

OpenStudy (anonymous):

35

OpenStudy (callmekiki):

Damn this is hard to explain.

OpenStudy (callmekiki):

Okay. 35. You know that. 1/7 is the first fraction. How many times does 7 go into 35?

OpenStudy (callmekiki):

5 times, so now multiply the numerator (1) by 5.

OpenStudy (callmekiki):

5 is the numerator and 35 is the denominator. 1/7 = 5/35

OpenStudy (callmekiki):

Now 3/5. 5 goes into 35 7 times. Multiply the numerator (3) by 7.

OpenStudy (callmekiki):

7 times 3 is 21

OpenStudy (callmekiki):

3/5 = 21/35

OpenStudy (callmekiki):

The new fractions are 5/35 and 21/35

OpenStudy (callmekiki):

Adding them will give you 26/35